how do yo solve
y=1-x
y=x-7

Answer by oscargut(2103)

y=1-x
y=x-7
Then
1-x = x-7
-2x = -8
x = 4
Answer: x = 4




If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks